# Speech Emotion Recognition with ASR Transcripts: A Comprehensive Study on Word Error Rate and Fusion Techniques

### v1

@@ 5.1 ASR Error Correction
We combine the transcripts from all used ASR models, forming N-best (N=11) hypotheses. Compared to using raw N-best hypotheses directly from a single ASR model, which have limited diversity, such a combination offers a more varied range of transcription results. Furthermore, this approach facilitates the comparison of SER performance with each ASR model, as depicted in Fig. 2.
Then, the ASR error correction consists of a two-stage process. First, we utilize the Alpaca prompt [31], which guides a Large Language Model (LLM) to generate the most likely transcript (i.e., 1-best hypothesis) from the N-best hypotheses and correct any grammatical or logical errors. An illustration of the Alpaca prompt is depicted on the right of Fig. 3. We employ InstructGPT, which has demonstrated effectiveness in LLM-based ASR error correction [32].
@@ 5.2 Modality-Gated Fusion
Based on the quality-enhanced transcripts, we adapt a trimodal incongruity-aware fusion method [7] to our bimodal source, considering potential distortion of original emotion in the spoken content due to ASR errors, which may introduce incongruence between emotions in audio and text.
Specifically, each modality is initially assigned an equal trainable weight based on its contribution to the final task. These weights are dynamically adjusted during training, with larger contributions resulting in higher values. The sum of the weights always equals 1, and they are updated in every training batch to ensure dynamic adaptation to any input batch size. Subsequently, the weights are multiplied with corresponding modality features to perform a weighted attention fusion. The pseudo-code is shown in Algorithm 1, where $H$ represents the hidden representation as input to the backbone SER model. $CrossAttn$ and $SelfAttn$ denote multihead attention, the former with the first argument as query and the second as key and value, and the latter with the argument as query, key, and value. The head number is set to eight.
@@
Input :  Audio modality $A$; Text modality $T$; Dynamic modality gating weight $W$ initialized with nn.Param(torch.tensor([1, 1]))
@@ 6      Sample $A$, $T$;
7      $W^{\prime}\leftarrow\texttt{softmax(}W\texttt{)}$;
8      $W_{A}\leftarrow W^{\prime}[0]$;
9      $W_{T}\leftarrow W^{\prime}[1]$;
@@
11             $A^{\prime}=W_{A}*A$;
12            $T^{\prime}=W_{T}*CrossAttn(A,T)$;
@@
14             $A^{\prime}=W_{A}*CrossAttn(T,A)$;
15            $T^{\prime}=W_{T}*T$;
@@ 18      $H=SelfAttn(Concat\left[A^{\prime};T^{\prime}\right])$;
Finally, we apply the same backbone SER model, keeping all other settings unchanged (e.g., the language and audio encoders) as in Section 4. We believe that this combination of dynamic modality gate and weighted attention can effectively mitigate the negative effects of ASR errors by consistently focusing on the most salient parts for SER.
@@ 5.3 Results and Discussion
The results are presented in Table 1. We only include MAE on CMU-MOSI and the average CCC on MSP-Podcast for brevity. Our approach further reduces the WER and improves SER performance, demonstrating its effectiveness. Although it does not surpass the best transcript in terms of MAE on CMU-MOSI, the values are very close. Moreover, consistent with the findings in [17], our results confirm that with proper ASR error-robust approaches, it is possible to exceed the SER performance of ground-truth text using ASR transcripts. Once again, this phenomenon likely occurs because certain misrecognized words are interpreted as contributing positively to the ground-truth emotions, rather than ASR errors being preferable in SER.
@@
best trans 12.31 74.46 18.69 0.8558 15.02 0.616
ground truth 0.00 74.19 0.00 0.8902 0.00 0.613
ours 10.12 76.66 17.00 0.8563 12.85 0.618
@@ 6 Conclusion
In this study, we conducted a benchmark of SER performance using ASR transcripts with varying WERs and explore mainstream fusion techniques to assess the impact of ASR performance on SER. Our findings revealed several novel insights: 1) SER can tolerate relatively low WER, especially in real-life speech scenarios; 2) bimodal SER with transcripts containing approximately 10% errors may not underperform those with ground-truth text, particularly with powerful Whisper models. However, further analysis is needed to understand what and where the ASR errors actually are, as well as the mechanisms underlying fusion techniques. Moreover, we proposed an ASR error-robust framework that integrates ASR error correction and modality-gated fusion. This framework demonstrated superior performance compared to baseline models in our experimental results and has the potential to find applications in various tasks where ASR transcription serves as the text source. In the near future, we plan to launch a challenge aimed at addressing ASR errors for robust multimodal SER.

### v2

@@ 5.1 ASR Error Correction
We combine the transcripts from all used ASR models, forming N-best (N=11) hypotheses. Compared to using raw N-best hypotheses directly from a single ASR model, which offer limited diversity, such a combination offers a more diverse range of transcription results. Furthermore, this approach facilitates the comparison of SER performance with each ASR model, as depicted in Fig. 2.
Then, the ASR error correction process consists of a two-stage process. First, we utilize the Alpaca prompt [32], which guides a Large Language Model (LLM) to generate the most likely transcript (i.e., 1-best hypothesis) from the N-best hypotheses and correct any grammatical or logical errors. An illustration of the Alpaca prompt is depicted on the right of Fig. 1. We employ InstructGPT, which has demonstrated effectiveness in LLM-based ASR error correction [33].
@@ 5.2 Modality-Gated Fusion
Based on the quality-enhanced transcripts, we adapt a trimodal incongruity-aware fusion method [7] to our bimodal inputs, considering potential distortion of original emotion in the spoken content due to ASR errors, which may introduce incongruence between emotions in audio and text.
Specifically, each modality is initially assigned an equal trainable weight. These weights dynamically adjust during training based on the respective contribution of each task, with larger contributions resulting in higher values. The sum of the weights always equals 1, and they are updated in every training batch to ensure dynamic adaptation to any input batch size. Subsequently, the weights are multiplied by corresponding modality features to perform weighted cross-attention fusion. The pseudo-code is shown in Algorithm 1, where $H$ represents the hidden representation as input to the backbone SER model. $CrossAttn$ and $SelfAttn$ denote multihead attention, the former with the first argument as query and the second as key and value, and the latter with the argument as query, key, and value. The head number is set to eight.
@@
Input :  Audio modality $A$; Text modality $T$; Dynamic modality gating weight $W^{1}$ initialized with nn.Param(torch.tensor([1, 1])); Feature concatenation weight $W^{2}$ initialized with nn.Param(torch.tensor([1, 1, 1]))
@@ 5      Sample $A$, $T$;
6      $W^{1}\leftarrow\texttt{Softmax(}W^{1}\texttt{)}$;
7      $W^{2}\leftarrow\texttt{Softmax(}W^{2}\texttt{)}$;
8      $W^{1}_{A}\leftarrow W^{1}[0]$, $W^{1}_{T}\leftarrow W^{1}[1]$;
9      $W^{2}_{A}\leftarrow W^{2}[0]$, $W^{2}_{T}\leftarrow W^{2}[1]$, $W^{2}_{AT}\leftarrow W^{2}[2]$,
@@
11             $A^{\prime}=W^{1}_{A}*A$;
12            $T^{\prime}=W^{1}_{T}*CrossAttn(A,T)$;
@@
14             $A^{\prime}=W^{1}_{A}*CrossAttn(T,A)$;
15            $T^{\prime}=W^{1}_{T}*T$;
@@ 19      $H^{\prime}=Concat\left[W^{2}_{A}*A,W^{2}_{T}*T,W^{2}_{AT}*H\right]$;
Finally, we apply the same backbone SER model, keeping all other settings unchanged (e.g., the text and audio encoders) as in Section 4. We believe that this combination of dynamic modality gate and weighted attention can effectively mitigate the negative impact of ASR errors by consistently focusing on the most salient parts in both inputs for SER.
@@ 5.3 Results and Discussion
The results are presented in Table 3. We include only the MAE on CMU-MOSI and the average CCC on MSP-Podcast for brevity. Our approach further reduces WER and improves SER performance, demonstrating its effectiveness. While it does not surpass the best transcript in terms of MAE on CMU-MOSI, the values are very close. Moreover, in line with the findings in [17], our results confirm that employing proper ASR error-robust approaches can exceed SER performance based on ground-truth text using ASR transcripts. This phenomenon likely occurs because certain misrecognized words are interpreted as contributing positively to the ground-truth emotions, rather than ASR errors being preferable in SER.
@@
best trans 12.31 74.66 18.69 0.8558 15.02 0.616
ground truth 0.00 74.32 0.00 0.8902 0.00 0.613
ours 10.12 76.66 17.00 0.8557 12.85 0.618
@@ 6 Conclusion
In this study, we conducted a benchmark of SER performance using ASR transcripts with varying WERs and explored mainstream fusion techniques to assess the impact of ASR performance on SER. Our findings revealed several novel insights: 1) SER can tolerate relatively low WERs, especially in real-life speech scenarios. 2) Bimodal SER with transcripts containing approximately 10% errors may not perform worse than those with ground-truth text, particularly with powerful Whisper models. However, further analysis is necessary to understand the nature and locations of ASR errors, as well as the mechanisms underlying fusion techniques. Moreover, we proposed an ASR error-robust framework that integrates ASR error correction and modality-gated fusion. This framework demonstrated superior performance compared to baseline models in our experimental results and holds potential for various tasks where ASR transcription serves as the text source. In the near future, we plan to launch a challenge aimed at addressing ASR errors for robust multimodal SER.